A 120V generator powers an electric circuit with 86.0 Ω of resistance .  

a. How much electric current flows in this circuit?   A


b. What is this circuit's electric power?   W


c. The generator runs for one hour. Calculate how much energy, in J, the generator will use.   J
Answer #1

a) using ohm's law

V=IR

I=V/R here V=120 V , R=86 ohm

I=120/86

I=1.395 A

b) P=I2/R

P=1.3952×86

P=167.35815

P=167.36 W.

c)

H=Pt

here P=167.36 W, t=1 hour =3600s

H=167.36×3600

H=602496 J Answer